i am 10 years old my father is 2 times older than my brother my mom is 10 years younger and my brother is 20 how old in my mother
Answer by ankor@dixie-net.com(22740)

I am 10 years old.
I = 10
:
My father is 2 times older than my brother.
f = 2b
:
My mom is 10 years younger and my brother is 20.
m = f - 10
b = 20
;
How old in my mother?
f = 2(20)
f = 40
and
m = 40 - 10
m = 30
:
It must be a half-brother, if mom is 30